OpenStudy (anonymous):

7(3y+2)-3=6y-2 What does y equal

OpenStudy (anonymous):

ok first distribute....

OpenStudy (anonymous):

21y+14-3=6y-2

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

7(3y+2)-3=6y-2 y= 13/15

OpenStudy (anonymous):

no direct answers

OpenStudy (anonymous):

15y+11=-2

OpenStudy (anonymous):

Y=-0.86666666667

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

did you still need help with any of those?

OpenStudy (xxx_gagnon_xxx):

hit me up i can help you any time of the day

OpenStudy (anonymous):

hahaha